Write an equation in slope-intercept form for the line that contains each pair of points below
(6,22) and (2,14)

Answer by rfer(16322)

m=14-22/2-6
m=-8/-4
m=2
-----------------------
y=mx+b solve for b
14=2*2+b
b=10
------------------
y=2x+10